OpenStudy (anonymous):

is the square root of 37 irrational or rational

OpenStudy (anonymous):

irrational

OpenStudy (anonymous):

thanks and y

OpenStudy (anonymous):

because its not a even number

OpenStudy (anonymous):

Actually, only perfect squares have rational square roots, i.e. sqrt(4) is rational since 2^2 = 4 sqrt(9) is rational since 3^2 = 9 sqrt(16) is rational since 4^2 = 16 etc...